# Decentralized Domain Generalization with Style Sharing: Formal Model and Convergence Analysis

### v3

@@ III-C Formal Modeling of Centralized DG
In this section, we provide the first formal models of the $\operatorname{MixStyle}$ [28] and $\operatorname{DSU}$ [29] algorithms for centralized DG. Note that the original papers and subsequent works do not present any such objective functions. This serves as a precursor to developing our StyleDDG algorithm in Sec. IV-B.
III-C1 $\operatorname{MixStyle}$ [28]
In $\operatorname{MixStyle}$, the style space of the CNN layers are explored through convex combinations of the existing styles, thereby exposing the model to styles beyond the existing ones in the training batch. For a given instance $x$ and target style statistics $\mu_{t}$ and $\sigma_{t}$, the $\operatorname{MixStyle}$ operation employs $\operatorname{AdaIN}$ from Eq. (1) as
$\underset{\lambda}{\operatorname{MixStyle}}(x,\mu_{t},\sigma_{t})=\operatorname{AdaIN}(x,\beta_{\lambda}(x,\mu_{t}),\gamma_{\lambda}(x,\sigma_{t})),$ (4)
@@ III-C1 $\operatorname{MixStyle}$ [28]
respectively, and $\lambda$ is the mixing coefficient. The original paper [28] applies $\operatorname{MixStyle}$ to all convolutional blocks except the last one for ResNet ($L=3$). They sample $\lambda$ for each layer from the Beta distribution $\mathbb{p}_{\lambda}=\operatorname{Beta}(0.1,0.1)$. Additionally, for each layer $\ell=1,...,L$, they assign a probability $p_{\ell}=0.5$ of applying $\operatorname{MixStyle}$ to layer $\ell$.
@@
where $\mathbb{p}_{D}$ denotes the probability distribution of data in the source domains and $\mathfrak{S}_{\mathcal{B}}$ is the set of all possible permutations of the set $\mathcal{B}=\{1,...,B\}$. In other words, [28] applies the $\operatorname{MixStyle}$ strategy to the outputs of the model’s layer $\ell$ with a probability $p_{\ell}$, so that the overall model is modified as
@@
in which $\hat{h}_{\theta_{\ell},I_{\ell},\lambda_{\ell}}=h_{\theta_{\ell}}$ if $\ell\notin\pi$. If $\ell\in\pi$, we define $x_{m}=[x_{k}]_{k\in I_{l}}^{T}$ as a permutated version of the mini-batch $x$ based on the shuffled indices $I_{\ell}$, and then the $\operatorname{MixStyle}$ update that occurs at that layer can be defined using Eq. (4) as
@@
Note that the $\operatorname{MixStyle}$ layer is taking the model outputs for mini-batch $x$ and the style statistics $\mu(\cdot)$ and $\sigma(\cdot)$ of the permuted mini-batch $x_{m}$ from Eq. (2) for feature augmentation.
III-C2 $\operatorname{DSU}$ [29]
@@
respectively, and $\epsilon_{\mu}$ and $\epsilon_{\sigma}$ follow a Normal distribution. Specifically, [29] applies $\operatorname{DSU}$ to all convolutional blocks alongside the first convolutional layer and the max pooling layer of a ResNet model ($L=6$). The exploration coefficients $\epsilon_{\mu,\ell},\epsilon_{\sigma,\ell}$ are sampled for each layer $\ell$ from distributions $\mathbb{p}_{\epsilon_{\mu}},\mathbb{p}_{\epsilon_{\sigma}}$ set as $\mathbb{p}_{\epsilon_{\mu}}=\mathbb{p}_{\epsilon_{\sigma}}=\mathcal{N}(0,1)$, i.e., unit Gaussian. As in $\operatorname{MixStyle}$, the probability of applying $\operatorname{DSU}$ to each layer $\ell=1,...,L$ is set to $p_{\ell}=0.5$.
@@
In essence, this method applies its $\operatorname{DSU}$ strategy to the outputs of the model’s layer $\ell$ with a probability $p_{\ell}$, so that the overall model is modified as
@@
in which $\dot{h}_{\theta_{\ell},\epsilon_{\mu,\ell},\epsilon_{\sigma,\ell}}=h_{\theta}$ if $\ell\notin\pi$. If $\ell\in\pi$, then the $\operatorname{DSU}$ update that occurs at that layer can be defined as
@@
where the $\operatorname{DSU}$ layer takes the model outputs for mini-batch $x$ and the higher-order style statistics $\Sigma_{\mu}(\cdot)$ and $\Sigma_{\sigma}(\cdot)$ of that batch as calculated in Eq. (3) to conduct feature augmentation.
The loss function formulations we derived in Eqs. (6) and (11) provide us with compact representations that will enable us to formally model our StyleDDG algorithm in Sec. IV. We observe that minimizing Eqs. (6) and (11) results in learning style-invariant model parameters for all different combinations of the layers/blocks in the CNN, by taking a convex combination of the styles (for $\operatorname{MixStyle}$) and adding Gaussian noise to the styles of each layer (for $\operatorname{DSU}$).
@@ IV-A1 Gradient Descent and Consensus Mechanism
In our decentralized network setup, each device $i\in\mathcal{M}$ maintains a local model $\theta_{i}^{(k)}$ at each iteration $k$ of the training process. A consensus mechanism is necessary to ensure $\lim_{k\to\infty}{\theta_{1}^{(k)}}=...=\lim_{k\to\infty}{\theta_{m}^{(k)}}$. To this end, we represent the update rule for each device $i\in\mathcal{M}$ at each iteration $k$ as in [41, 15]:
$\theta_{i}^{(k+1)}=\theta_{i}^{(k)}+\sum_{j\in\mathcal{N}_{i}}{p_{ij}\left(\theta_{j}^{(k)}-\theta_{i}^{(k)}\right)}-\alpha^{(k)}g_{i}^{(k)},$ (14)
where $p_{ij}$ is a mixing coefficient used for the model parameters of device pair $(i,j)$ with a D2D link. Here, we employ the commonly used Metropolis-Hastings weights $p_{ij}=\min\{1/(1+|\mathcal{N}_{i}|),1/(1+|\mathcal{N}_{j}|)\}$ [42]. Furthermore, $g_{i}^{(k)}=\nabla{F}_{i}(\theta_{i}^{(k)},[\psi_{j}^{(k)}]_{j\in\mathcal{N}_{i}})$ is the style-based gradient of device $i\in\mathcal{M}$ at iteration $k$ obtained based on the local loss function, which will be elaborated in Sec. IV-A3. It has been well-established how the convergence speed of DFL in (14) decreases with the spectral radius $\rho$ of the mixing matrix on the D2D graph $\mathcal{G}$ [41, 43]; a core contribution of our work is expanding such results to style sharing in DG.
@@ IV-A2 Style Statistics
Let $X\in\mathbb{R}^{m\times B}$ be the concatenation of all local mini-batches of data from the devices, i.e., $X=\begin{bmatrix}x_{1}&...&x_{m}\end{bmatrix}^{T}$. We denote the set of styles from all devices obtained using model parameters $\theta$ and combined batch $X$ for all devices as $\Psi_{\theta}(X)=\begin{bmatrix}\psi_{\theta}(x_{1})&\cdots&\psi_{\theta}(x_{m})\end{bmatrix}^{T}$. The style tensor that each device $i\in\mathcal{M}$ calculates and shares with its neighbors is defined as $\psi_{\theta}(x_{i})=\begin{bmatrix}\psi_{\theta,1}(x_{i})&\cdots&\psi_{\theta,L}(x_{i})\end{bmatrix}^{T}$, which is obtained for different CNN layers/blocks. Each individual style in this tensor consists of four style statistics at each layer/block $\psi_{\theta,\ell}(x_{i})=\begin{bmatrix}\bar{\mu}_{\theta,\ell}(x_{i})&\bar{\sigma}_{\theta,\ell}(x_{i})&\Sigma_{\mu,\theta,\ell}^{2}(x_{i})&\Sigma_{\sigma,\theta,\ell}^{2}(x_{i})\end{bmatrix}^{T}$:
$\begin{gathered}\bar{\mu}_{\theta,\ell}(x_{i})_{c}=\frac{1}{B}\sum_{b=1}^{B}{\mu(h_{\theta_{\ell}}(x_{i}))_{b,c}},\\ \bar{\sigma}_{\theta,\ell}(x_{i})_{c}=\frac{1}{B}\sum_{b=1}^{B}{\sigma(h_{\theta_{\ell}}(x_{i}))_{b,c}},\\ \Sigma_{\mu,\theta,\ell}^{2}(x_{i})_{c}=\Sigma_{\mu}^{2}(h_{\theta_{\ell}}(x_{i}))_{c},\;\Sigma_{\sigma,\theta,\ell}^{2}(x_{i})_{c}=\Sigma_{\sigma}^{2}(h_{\theta_{\ell}}(x_{i}))_{c},\end{gathered}\hskip-28.45274pt$ (15)
@@
Therefore, the dimensions of the style vectors will be $\psi_{\theta_{\ell}}(x_{i})\in\mathbb{R}^{4\times C}$ for each layer, $\psi_{\theta}(x_{i})\in\mathbb{R}^{L\times 4\times C}$ for each device and $\Psi_{\theta}(X)\in\mathbb{R}^{m\times L\times 4\times C}$ for the full network. Fig. 2 gives a detailed illustration of these style vectors and statistics.
@@
(Communication overhead) The added communication overhead for sharing the style statistics $\psi_{\theta}(x_{i})\in\mathbb{R}^{L\times 4\times C}$ is negligible relative to the model parameters $\theta\in\mathbb{R}^{n}$. Take the ResNet18 CNN model as an example, with around $n=11.69$ million parameters. The number of parameters in the full style statistics vector in StyleDDG will be $1792$ (since there are four equidimensional style statistics at each layer, and they are calculated in the outputs of the first three convolutional blocks with channel numbers $64$, $128$ and $256$, respectively). As a result, the added communication overhead is $0.015\%$, which is negligible.
@@ IV-A3 Style Exploration Strategy
In StyleDDG, the goal of each device is to communicate style information with its neighbors, so that they can collectively train domain-invariant models. Towards this, we apply three functions to the outputs of particular layers of a CNN, namely, $\operatorname{StyleShift}$, $\operatorname{StyleExplore}$ and $\operatorname{MixStyle}$. In the following subsections, we explain the details of each of them.
@@
$\mathbf{1}_{B_{s}}$ is an all-ones vector of size $B_{s}$, and $B$ is the batch size. Recall that the $\operatorname{AdaIN}$ function was defined in Eq. (1). The shifting and scaling hyperparameters are the same ones used for $\operatorname{DSU}$ as defined in Eq. (10).
StyleExplore. After shifting the styles of $B_{s}$ points, we perform a subsequent step called style exploration, to further expand the style space each device is exposed to during training. To this end, we first calculate intra-device average of style statistics, i.e., $\mathbb{E}_{b}[\mu(x)]$ and $\mathbb{E}_{b}[\sigma(x)]$. Then, we use those average points to extrapolate $B_{e}\in\mathbb{Z},B_{e}\in[0,B]$ of $\operatorname{StyleShift}$ outputs outwards of the style space, as illustrated in Fig. 1-(b). To mathematically model the points being extrapolated, we let the $\operatorname{StyleExplore}$ function take a binary vector $I_{e}$ with $1^{T}I_{e}=B_{e}$ as input, and shift those points in $x$ which correspond to $1$s in $I_{e}$. In other words, for a general given vector $x$, we choose $B_{e}$ of it $x_{e}=[x_{k}]_{k\in\{1,...,B\};(I_{e})_{k}=1}^{T}$ to apply style extrapolation on and keep the rest $x_{e}^{c}=[x_{k}]_{k\in\{1,...,B\};(I_{e})_{k}=1}^{T}$ the same. We finally concatenate them to get
@@
where $\alpha$ is an exploration hyperparameter set to $\alpha=3$ as in [13]. Different than [13], we do not do any oversampling in this layer to avoid doubling the computation overhead of our methodology. This is a very critical modification to make, since other well-known DG methods such as [28, 29, 32] all operate with the rationale that style exploration layers should be plug-and-play components of existing models.
MixStyle. In the end, we use another set of indices $I_{m}$ (a 0-1 vector similar to $I_{e}$) to obtain a permutation of $\mu_{e}$ and $\sigma_{e}$ and use them to do $\operatorname{MixStyle}$. The $\operatorname{MixStyle}$ function was given in Eq. (4). This step of our method is also illustrated in Fig. 1-(c). The final function can be written as
@@
1
[…]
3while $k\leq K$ do
4  forall $i\in\mathcal{M}$ do
5     sample mini-batch $(x_{i}^{(k)},y_{i}^{(k)})\in\mathcal{D}_{i}$;
6    calculate $\psi_{i}^{(k)}=\psi_{\theta_{i}^{(k)}}(x_{i}^{(k)})$ based on Eq. (15); share $\theta_{i}^{(k)}$ and $\psi_{i}^{(k)}$ with neighbors $\mathcal{N}_{i}$;
7
8 forall $i\in\mathcal{M}$ do
9
10    use the same mini-batch $(x_{i}^{(k)},y_{i}^{(k)})$ from Line 1;
11    $g_{i}^{(k)}\leftarrow\nabla{F}_{i}(\theta_{i}^{(k)},[\psi_{j}^{(k)}]_{j\in\mathcal{N}_{i}})$ (with respect to functions explained in Sec. IV-A3);
12    $\text{aggr}_{i}^{(k)}\leftarrow 0$;
13    forall $j\in\mathcal{E}_{i}$ do
14        $p_{ij}\leftarrow 1/(1+\max{\{|\mathcal{N}_{i}|,|\mathcal{N}_{j}|\}})$;
15       $\text{aggr}_{i}^{(k)}\leftarrow\text{aggr}_{i}^{(k)}+p_{ij}\left(\mathbf{\theta}_{j}^{(k)}-\mathbf{\theta}_{i}^{(k)}\right)$;
16
17
18 for $i\in\mathcal{M}$ do
19     $\mathbf{\theta}_{i}^{(k+1)}\leftarrow\mathbf{\theta}_{i}^{(k)}+\text{aggr}_{i}^{(k)}-\alpha^{(k)}g_{i}^{(k)}$;
20 $k\leftarrow k+1$;
@@
Novelty Summary. In a nutshell, the novelty of our algorithm is two-fold. First, StyleDDG is, to our knowledge, the first style-based algorithm for achieving DG over fully decentralized D2D networks. The consensus-based decentralized update rule in Eq. (14) is used with modified gradients $\nabla{F}_{i}(\theta_{i}^{(k)},[\psi_{j}^{(k)}]_{j\in\mathcal{N}_{i}})$ to achieve DG. Second, our style-based approach utilizes style information sharing only between the one-hop neighbors of a device $i\in\mathcal{M}$, and is still able to train style-invariant ML models, as we will see next.

### v4

@@ III-C Formal Modeling of Centralized DG
In this section, we provide the first formal models of the $\operatorname{MixStyle}$ [zhou2021domain] and $\operatorname{DSU}$ [li2022uncertainty] algorithms for centralized DG. Note that the original papers and subsequent works do not present any such objective functions. This serves as a precursor to developing our StyleDDG algorithm in Sec. IV-B.
III-C1 $\operatorname{MixStyle}$ [zhou2021domain]
In $\operatorname{MixStyle}$, the style space of the CNN layers is explored through convex combinations of the existing styles, thus exposing the model to styles beyond the existing ones in the training batch. For a given instance $x$ and target style statistics $\mu_{t}$ and $\sigma_{t}$, the $\operatorname{MixStyle}$ operation employs $\operatorname{AdaIN}$ from Eq. (1) as
$\underset{\lambda}{\operatorname{MixStyle}}(x,\mu_{t},\sigma_{t})=\operatorname{AdaIN}(x,\beta_{\lambda}(x,\mu_{t}),\gamma_{\lambda}(x,\sigma_{t})),\vskip-7.22743pt$ (4)
@@ III-C1 $\operatorname{MixStyle}$ [zhou2021domain]
respectively, and $\lambda$ is the mixing coefficient. The original paper [zhou2021domain] applies $\operatorname{MixStyle}$ to all convolutional blocks except the last one for ResNet ($L=3$), and samples $\lambda$ for each layer using the Beta distribution $\mathbb{p}_{\lambda}=\operatorname{Beta}(0.1,0.1)$. Additionally, for each layer $\ell=1,...,L$, they assign a probability $p_{\ell}=0.5$ of applying $\operatorname{MixStyle}$ to layer $\ell$.
@@
where $\mathbb{p}_{D}$ denotes the probability distribution of the data in the source domains and $\mathfrak{S}_{\mathcal{B}}$ is the set of all possible permutations of the set $\mathcal{B}=\{1,...,B\}$. In other words, [zhou2021domain] applies the $\operatorname{MixStyle}$ strategy to the output of the model’s layer $\ell$ with a probability $p_{\ell}$, so that the overall model is modified as
@@
in which $\hat{h}_{\theta_{\ell},I_{\ell},\lambda_{\ell}}=h_{\theta_{\ell}}$ if $\ell\notin\pi$. If $\ell\in\pi$, we define $x_{m}=[x_{k}]_{k\in I_{l}}^{T}$ as a permutated version of the mini-batch $x$ based on the shuffled indices $I_{\ell}$, and then the $\operatorname{MixStyle}$ update that occurs in that layer can be defined using Eq. (4) as
@@
Note that the $\operatorname{MixStyle}$ layer is taking the model output for mini-batch $x$ and the style statistics $\mu(\cdot)$ and $\sigma(\cdot)$ of the permuted mini-batch $x_{m}$ from Eq. (2) for feature augmentation.
III-C2 $\operatorname{DSU}$ [li2022uncertainty]
@@
respectively, and $\epsilon_{\mu}$ and $\epsilon_{\sigma}$ follow a Normal distribution. Specifically, [li2022uncertainty] applies $\operatorname{DSU}$ to all convolutional blocks along with the first convolutional layer and the max-pooling layer of a ResNet model ($L=6$). The exploration coefficients $\epsilon_{\mu,\ell},\epsilon_{\sigma,\ell}$ are sampled for each layer $\ell$ from distributions $\mathbb{p}_{\epsilon_{\mu}},\mathbb{p}_{\epsilon_{\sigma}}$ set as $\mathbb{p}_{\epsilon_{\mu}}=\mathbb{p}_{\epsilon_{\sigma}}=\mathcal{N}(0,1)$, i.e., unit Gaussian distribution. As in $\operatorname{MixStyle}$, the probability of applying $\operatorname{DSU}$ to each layer $\ell=1,...,L$ is set to $p_{\ell}=0.5$.
@@
In essence, this method applies its $\operatorname{DSU}$ strategy to the output of the model’s layer $\ell$ with probability $p_{\ell}$, so that the overall model is modified as
@@
in which $\dot{h}_{\theta_{\ell},\epsilon_{\mu,\ell},\epsilon_{\sigma,\ell}}=h_{\theta}$ if $\ell\notin\pi$. If $\ell\in\pi$, then the $\operatorname{DSU}$ update that occurs in that layer can be defined as
@@
where the $\operatorname{DSU}$ layer takes the model outputs for the mini-batch $x$ and the higher-order style statistics $\Sigma_{\mu}(\cdot)$ and $\Sigma_{\sigma}(\cdot)$ of that batch as given in Eq. (3) to perform feature augmentation.
The loss function formulations we derived in Eqs. (6) and (11) provide us with compact representations that will enable us to formally model our StyleDDG algorithm in Sec. IV. We observe that minimizing Eqs. (6) and (11) results in learning style-invariant model parameters for all different combinations of layers/blocks in the CNN, by taking a convex combination of the styles (for $\operatorname{MixStyle}$) and adding Gaussian noise to the styles of each layer (for $\operatorname{DSU}$).
@@ IV-A1 Gradient Descent and Consensus Mechanism
In our decentralized network setup, each device $i\in\mathcal{M}$ maintains a local model $\theta_{i}^{(k)}$ at each iteration $k$ of the training process. A consensus mechanism is necessary to ensure $\lim_{k\to\infty}{\theta_{1}^{(k)}}=...=\lim_{k\to\infty}{\theta_{m}^{(k)}}$. To this end, we represent the update rule for each device $i\in\mathcal{M}$ in each iteration $k$ as in [zehtabi2024decentralized, chen2024taming]:
$\theta_{i}^{(k+1)}=\theta_{i}^{(k)}+\sum_{j\in\mathcal{N}_{i}}{p_{ij}\left(\theta_{j}^{(k)}-\theta_{i}^{(k)}\right)}-\alpha^{(k)}g_{i}^{(k)},\vskip-7.22743pt$ (14)
where $p_{ij}$ is a mixing coefficient used for the model parameters of the device pair $(i,j)$ with a D2D link. Here, we employ the commonly used Metropolis-Hastings weights $p_{ij}=\min\{1/(1+|\mathcal{N}_{i}|),1/(1+|\mathcal{N}_{j}|)\}$ [xiao2004fast]. Furthermore, $g_{i}^{(k)}=\nabla{F}_{i}(\theta_{i}^{(k)},[\psi_{j}^{(k)}]_{j\in\mathcal{N}_{i}})$ is the style-based gradient of device $i\in\mathcal{M}$ at iteration $k$ obtained based on the local loss function, which will be elaborated in Sec. IV-A3. It has been well-established how the convergence speed of DFL in (14) decreases with the spectral radius $\rho$ of the mixing matrix on the D2D graph $\mathcal{G}$ [zehtabi2024decentralized, koloskova2020unified]; a main contribution of our work is expanding such results to style sharing in DG.
@@ IV-A2 Style Statistics
Let $X\in\mathbb{R}^{m\times B}$ be the concatenation of all local mini-batches of data from the devices, i.e., $X=\begin{bmatrix}x_{1}&...&x_{m}\end{bmatrix}^{T}$. We denote the set of styles from all the devices obtained using model parameters $\theta$ and the combined batch $X$ for all devices as $\Psi_{\theta}(X)=\begin{bmatrix}\psi_{\theta}(x_{1})&\cdots&\psi_{\theta}(x_{m})\end{bmatrix}^{T}$. The style tensor that each device $i\in\mathcal{M}$ calculates and shares with its neighbors is defined as $\psi_{\theta}(x_{i})=\begin{bmatrix}\psi_{\theta,1}(x_{i})&\cdots&\psi_{\theta,L}(x_{i})\end{bmatrix}^{T}$, which is obtained for different CNN layers/blocks. Each individual style in this tensor consists of four style statistics at each layer/block $\psi_{\theta,\ell}(x_{i})=\begin{bmatrix}\bar{\mu}_{\theta,\ell}(x_{i})&\bar{\sigma}_{\theta,\ell}(x_{i})&\Sigma_{\mu,\theta,\ell}^{2}(x_{i})&\Sigma_{\sigma,\theta,\ell}^{2}(x_{i})\end{bmatrix}^{T}$:
$\vskip-7.22743pt\begin{gathered}\bar{\mu}_{\theta,\ell}(x_{i})_{c}=\frac{1}{B}\sum_{b=1}^{B}{\mu(h_{\theta_{\ell}}(x_{i}))_{b,c}},\\ \bar{\sigma}_{\theta,\ell}(x_{i})_{c}=\frac{1}{B}\sum_{b=1}^{B}{\sigma(h_{\theta_{\ell}}(x_{i}))_{b,c}},\\ \Sigma_{\mu,\theta,\ell}^{2}(x_{i})_{c}=\Sigma_{\mu}^{2}(h_{\theta_{\ell}}(x_{i}))_{c},\;\Sigma_{\sigma,\theta,\ell}^{2}(x_{i})_{c}=\Sigma_{\sigma}^{2}(h_{\theta_{\ell}}(x_{i}))_{c},\end{gathered}\hskip-28.45274pt$ (15)
@@
Therefore, the dimensions of the style vectors will be $\psi_{\theta_{\ell}}(x_{i})\in\mathbb{R}^{4\times C}$ for each layer, $\psi_{\theta}(x_{i})\in\mathbb{R}^{L\times 4\times C}$ for each device, and $\Psi_{\theta}(X)\in\mathbb{R}^{m\times L\times 4\times C}$ for the whole network. Fig. 2 illustrates these style vectors and statistics.
@@
(Communication overhead) The added communication overhead for sharing style statistics $\psi_{\theta}(x_{i})\in\mathbb{R}^{L\times 4\times C}$ is negligible relative to the model parameters $\theta\in\mathbb{R}^{n}$. Take ResNet18 as an example, with around $n=11.69$ million parameters. The length of the full style statistics vector in StyleDDG is just $1792$ (four equidimensional style statistics in each layer, calculated in the outputs of the first three convolutional blocks with $64$, $128$ and $256$ channels, respectively), giving added communication overhead of just $0.015\%$.
@@ IV-A3 Style Exploration Strategy
In StyleDDG, the goal of each device is to communicate style information with its neighbors so that they can collectively train domain-invariant models. Towards this, we apply three functions to the outputs of particular layers of a CNN, namely $\operatorname{StyleShift}$, $\operatorname{StyleExplore}$ and $\operatorname{MixStyle}$. In the following subsections, we explain the details of each of them.
@@
$\mathbf{1}_{B_{s}}$ is an all-ones vector of size $B_{s}$, and $B$ is the batch size. Recall that the function $\operatorname{AdaIN}$ was defined in Eq. (1). The shifting and scaling hyperparameters are the same ones used for $\operatorname{DSU}$ as defined in Eq. (10).
StyleExplore. After shifting the styles of $B_{s}$ points, we perform another step called style exploration to further expand the style space that each device is exposed to during training. To this end, we first calculate the intra-device average of style statistics, i.e., $\mathbb{E}_{b}[\mu(x)]$ and $\mathbb{E}_{b}[\sigma(x)]$. Then, we use those average points to extrapolate $B_{e}\in\mathbb{Z},B_{e}\in[0,B]$ of the outputs of $\operatorname{StyleShift}$ outward of the style space, as illustrated in Fig. 1-(b). To mathematically model the points that are extrapolated, we let the function $\operatorname{StyleExplore}$ take a binary vector $I_{e}$ with $1^{T}I_{e}=B_{e}$ as input and shift those points in $x$ which correspond to $1$s in $I_{e}$. In other words, for any given vector $x$, we choose $B_{e}$ of it $x_{e}=[x_{k}]_{k\in\{1,...,B\};(I_{e})_{k}=1}^{T}$ to apply style extrapolation on and keep the rest $x_{e}^{c}=[x_{k}]_{k\in\{1,...,B\};(I_{e})_{k}=1}^{T}$ the same. We finally concatenate them to get
@@
where $\alpha$ is an exploration hyperparameter set to $\alpha=3$ as in [park2024stablefdg]. Unlike [park2024stablefdg], we do not perform any oversampling in this layer to avoid doubling the computational overhead of our methodology. This is a very critical modification to make, since other well-known DG methods such as [zhou2021domain, li2022uncertainty, lifedbn] all operate with the rationale that style exploration layers should be plug-and-play components of existing models.
MixStyle. As a final step, we use another set of indices $I_{m}$ (a 0-11 vector similar to $I_{e}$) to obtain a permutation of $\mu_{e}$ and $\sigma_{e}$ and use them to do $\operatorname{MixStyle}$, with function $\operatorname{MixStyle}$ defined in Eq. (4). This step of our method is also illustrated in Fig. 1-(c). The final function can be written as
@@
$k\leftarrow 0$, Initialize $\theta^{(0)}$, ${\{\mathbf{\theta}_{i}^{(0)}\leftarrow\theta^{(0)}\}}_{i\in\mathcal{M}}$;
while $k\leq K$ do
  forall $i\in\mathcal{M}$ do
     sample mini-batch $(x_{i}^{(k)},y_{i}^{(k)})\in\mathcal{D}_{i}$;
    calculate $\psi_{i}^{(k)}=\psi_{\theta_{i}^{(k)}}(x_{i}^{(k)})$ using Eq. (15); share $\theta_{i}^{(k)}$ and $\psi_{i}^{(k)}$ with neighbors $\mathcal{N}_{i}$;
  end forall
 forall $i\in\mathcal{M}$ do
    use the same mini-batch $(x_{i}^{(k)},y_{i}^{(k)})$ from Line 1;
    $g_{i}^{(k)}\leftarrow\nabla{F}_{i}(\theta_{i}^{(k)},[\psi_{j}^{(k)}]_{j\in\mathcal{N}_{i}})$ (with respect to functions explained in Sec. IV-A3);
    $\text{aggr}_{i}^{(k)}\leftarrow 0$;
    forall $j\in\mathcal{E}_{i}$ do
        $p_{ij}\leftarrow 1/(1+\max{\{|\mathcal{N}_{i}|,|\mathcal{N}_{j}|\}})$;
       $\text{aggr}_{i}^{(k)}\leftarrow\text{aggr}_{i}^{(k)}+p_{ij}\left(\mathbf{\theta}_{j}^{(k)}-\mathbf{\theta}_{i}^{(k)}\right)$;
     end forall
  end forall
 for $i\in\mathcal{M}$ do
     $\mathbf{\theta}_{i}^{(k+1)}\leftarrow\mathbf{\theta}_{i}^{(k)}+\text{aggr}_{i}^{(k)}-\alpha^{(k)}g_{i}^{(k)}$;
  end for
 $k\leftarrow k+1$;
end while
@@
Novelty Summary. In a nutshell, the novelty of our algorithm is two-fold. First, StyleDDG is, to our knowledge, the first style-based algorithm for achieving DG over fully decentralized D2D networks. The consensus-based decentralized update rule in Eq. (14) is used with modified gradients $\nabla{F}_{i}(\theta_{i}^{(k)},[\psi_{j}^{(k)}]_{j\in\mathcal{N}_{i}})$ to achieve DG. Second, our style-based approach utilizes the sharing of style information only between the one-hop neighbors of a device $i\in\mathcal{M}$, and is still able to train style-invariant ML models, as we will see next.